OpenStudy (awesome781):

g(x)=X^2-1 Find the x and y intercept

OpenStudy (anonymous):

do you mean g(X)=x^2-1 ???

OpenStudy (awesome781):

oh yes sorry

OpenStudy (solomonzelman):

\(\large\color{blue}{ \rm g(x)=x^2-1 }\) \(\large\color{blue}{ \rm HINT:~~~~~~~~~~~~~~~x^2-1=(x-1)(x+1) }\)

OpenStudy (anonymous):

1 and -1

OpenStudy (awesome781):

oh so i had to factor it to get my x and y?

OpenStudy (anonymous):

you would factor it and then set each factor, (x-1) and (x+1) equal to zero to find your x-intercept

OpenStudy (awesome781):

oh ok thanks for the help :)

OpenStudy (solomonzelman):

Have you learned about imaginary numbers (such as ` i ` ) ?

OpenStudy (awesome781):

i have a question is this question same thing when written as (x^2-1) in parenthesis?

OpenStudy (solomonzelman):

Well, when you write ` 3-1 ` , and ` (3-1) ` the meaning is the same. THe difference would be between \(\large\color{blue}{ \rm f(x)=x^2-1 }\) and \(\large\color{blue}{ \rm f(x)=(x-1)^2 }\)

OpenStudy (awesome781):

yes but my teacher said we are not messing with imaginary numbers yet

OpenStudy (solomonzelman):

cool. I am just saying that if you had a plus instead of minus, in \(\large\color{blue}{ \rm f(x)=x^2-1 }\) you would need imag. .....

OpenStudy (awesome781):

oh ok thanks again :)

OpenStudy (solomonzelman):

Anytime :)